OpenStudy (anonymous):

Write the slope-intercept equation for the line through (3, 7) and (9, 3). A. y = -3/2 x + 7 B. y = -2/3 x + 9 C. y = 3/2 x + 3

OpenStudy (anonymous):

plug in those points and see which equation hold true... or you can do it the long way...

OpenStudy (anonymous):

so to find m you have to use slope formula which is M= (y2-y1)/(x2-x1) so it turns out to be -2/3 then you plug in one set of point to solve for b. ex: 7=3(-2/3) +b B= 9 so the final answer is y = -2/3x + 9